OpenStudy (anonymous):

If 6 times a certain number is added to 8, the result is 32. Which of the following equations could be used to solve the problem? 6x = 32 6(x + 8) = 32 6x = 8 + 32 6x + 8 = 32

OpenStudy (♪chibiterasu):

The last one. 6x is 6 times the certain number and then added to 8.

OpenStudy (anonymous):

so its the lat one?

OpenStudy (texaschic101):

I agree

OpenStudy (♪chibiterasu):

Yes, the last one.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (♪chibiterasu):

The solution would be x = 4 6(4) + 8 24 + 8 32

OpenStudy (anonymous):

ohhh i c

OpenStudy (anonymous):

It s D .!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

In the solution of the equation 5 - 3x = 2x + 9, 3x is added to the equation first. Which of the following should be done next? add -9 add -5 add -2x

OpenStudy (texaschic101):

5 - 3x = 2x + 9 (add 3x) 5 = 5x + 9 (subtract 9 or add -9) - 4 = 5x x = -4/5 you would add -9 to both sides